An 8 meter wide concrete road pavement 500 meters long is desired to be constructed over a well-compacted gravel road, together with the necessary concrete curbs and gutters on both sides. In order to put the subgrade on an even level grade, a 500 cu.m. of sand filling is necessary, over which the 25-cm concrete pavement will be placed. How much will the project cot, allowing for 15% contingency, and assuming the following data:

(a)   Sandfill, including rolling and watering – ₱10.00 per cu.m.

(b)   Concrete curb and gutter (labor and materias) – ₱12.00 per linear meter.

(c)    Concrete pavement, 25-cm thick (labor and materials), including curing – ₱22.00 per sq.m.


1
Expert's answer
2021-02-15T18:23:23-0500
"1.15\times (500\times10)=5,750"


"1.15\times (1,000\times12)=13,800"


"1.15\times (8\times500\times22)=101,200"

Total:


"5,750+13,800+101,200=120,750"